1/1

Manfred Hausmann,

Home sickness

translated by N. N. 68

You also come from my home town, You must know all the home town news. At dawn, before the silken window, Is it too cold for plum blossom to show?

Translation of the German poem “Heimweh” by Manfred Hausmann (∗1898-09-10 – †1986-08-06) into English by N. N. 68.

Arg-925-1851 (2013-08-30 21:18:45)

Tiu ĉi angligo troviĝas en http: // www. chinese-poems. com/ plum. html . Tie estas indikita ankaŭ la origina ĉina-lingva versio de la poemo en ĉinaj literoj kaj en la latins pinjin-transskribo.

http://www.poezio.net/version?poem-id=925&version-id=1851